# EDGE-INFERNO: Simulating every observable star in faint dwarf galaxies and their consequences for resolved-star photometric surveys

### v1

@@ 1 Introduction
In recent years, numerical models of galaxies have unlocked the ability to account for individual stars within a whole galaxy (e.g. Emerick et al. 2018; Andersson et al. 2020, 2021, 2023; Lahén et al. 2020, 2023, 2024; Calura et al. 2022; Gutcke et al. 2021, 2022a, 2022b; Hislop et al. 2022; Steinwandel et al. 2023, 2024; Deng et al. 2024), as opposed to aggregating stars into a single stellar population particle. By explicitly drawing from the initial mass function (IMF) at stellar birth and modeling the stellar evolution tracks of each star, these models provide a much more detailed and robust account of how, when, and where feedback and metals are deposited into the interstellar medium (ISM). Furthermore, this novel model approach opens up the door for incorporating aspects of collisional dynamics present in star clusters that potentially affect both the ISM, galactic winds, and circumgalactic medium (e.g., runaway stars, Andersson et al., 2020, 2023; Steinwandel et al., 2023).
@@
In the rare cases when individual-star models have included a full cosmological environment (e.g. Gutcke et al. 2022a, b; Calura et al. 2022), only massive (and thus less numerous) stars ($\gtrapprox 4\,\,{\rm M}_{\odot}$) are individually sampled to ease the computational load. But slow-moving stellar mass loss from low-mass evolved stars has characteristic velocities ($\approx 10\,\rm km\,s^{-1}$; Höfner & Olofsson 2018) comparable to the depth of gravitational potential wells of faint dwarf galaxies. As a result, winds from e.g. asymptotic giant branch stars (AGBs) likely play a role in chemical enrichment and in regulating star formation activity (e.g. Rey et al. 2020). Furthermore, the lowest mass galaxies has star formation quenched during the epoch of reionization, therefore massive stars are long gone by $z=0$. Instead dwarf galaxy properties are characterized through the imaging and spectroscopy of the light from their old, evolved low-mass stars, motivating a complete account of the stellar population.
In this paper, we combine all these advances for the first time to take a major step toward interpreting resolved-star data from dwarf galaxies. We perform high-resolution cosmological simulations within the EDGE [^1] Engineering Dwarfs at Galaxy formation’s Edge framework, combining them with the star-by-star INFERNO [^2] INdividual stars with Feedback, Enrichment, and Realistic Natal mOtions model (Andersson et al., 2023). With this, we achieve a high numerical resolution ($3\,{\rm pc}$ at $z=0$) that allows us to capture the evolution of supernova (SN) explosions self-consistently over the full cosmological formation of the dwarf, while individually sampling all stars down to $0.5\,\,{\rm M}_{\odot}$ (main sequence lifetime $>14\,\rm Gyr$). This ensures that all major sources of stellar feedback are accounted for star by star, while self-consistently populating the observable spatial and color-magnitude distribution at $z=0$ by a considerable margin.
@@ 2 Star-by-star cosmological simulations
Star formation follows Agertz et al. (2020), assuming stars are born in quantities of $500\,{\rm M}_{\odot}$ in cells with gas density $>300\,{\rm cm}^{-3}$ and temperature $<10^{3}\,{\rm K}$. Spawning stars in groups of $500\,{\rm M}_{\odot}$ minimizes issues related to incomplete sampling of the IMF (Smith, 2021). In addition to the gas velocity, stars are stirred with a velocity dispersion of $0.01\,{\rm km\,s}^{-1}$ at birth (Andersson et al., 2023).
@@
Stellar feedback in our models includes SNe (type II and Ia) and stellar winds (from OB type and AGB stars) implemented as in Andersson et al. (2023). Radiation is not included in our suite of simulations (or in the simulations from Rey et al. 2019 we compare to) and is expected to further regulate star formation, systematically decreasing the total stellar masses of our galaxies (Agertz et al. 2020, Rey et al. in prep). We read the chemical yields and mass-loss of each star as a function of mass and metallicity using tabulated values from NuGrid (Pignatari et al., 2016; Ritter et al., 2018) for SNe type II and stellar winds, and from Seitenzahl et al. (2013) for SNe Ia. In all cases, we calculate the main sequence lifetime of a star following the fitting function from Raiteri et al. (1996). SNe Ia follow the time-delay distribution for field galaxies from Maoz & Graur (2017) assuming a rate of $I_{\rm Ia}=2.6\times 10^{-13}\,{\rm yr}^{-1}\,{\rm M}_{\odot}^{-1}$ starting $38\,{\rm Myr}$s after birth.
As in Andersson et al. (2023), massive stars ($8\leq m_{\star}\leq 60\,{\rm M}_{\odot}$) inject fast ($1000\,{\rm km\,s}^{-1}$) winds while on the main sequence and end their life with core-collapse directly injecting material and $10^{51}\,{\rm erg}$ if $m_{\star}\leq 30\,{\rm M}_{\odot}$ (assuming direct collapse to black hole otherwise). The high resolution of our simulations allows this direct injection of SNe energy and ensures accurate capture of the final momentum (Agertz et al. 2020).
@@ 3 Stellar mass growth
Compared with the traditional stellar-particle approach (dotted; Rey et al. 2019), swapping to individual-star feedback systematically increases $M_{\star}$ at all redshifts, by at most a factor of 2. Such variations are well within theoretical uncertainties at this mass scale (see, e.g., the discussion in Agertz et al. 2020, and review in Sales et al. 2022) which should be expected given the extensive updates to the chemical enrichment channels (and thus gas cooling) and updates to the stellar feedback implementation [^3] Even if the total feedback budget and injection mechanisms are fixed, the location and clustering of individual stars can be significantly different and affect galactic outflows (e.g. Andersson et al. 2020, 2023; Smith 2021)..
@@ 4.1 Color-magnitude diagrams
All our dwarf galaxies have diffuse light profiles, with hierarchical merging leading to distant bound members up to $\geq 5\,r_{h}$ similar to those reported in recent observations (e.g., Chiti et al. 2021; Sestito et al. 2023). As we will see, these multi-component spatial distributions can bias the recovery of faint dwarf structural parameters, particularly when few stars are bright enough to be observed. Such outer stellar halo for faint dwarf galaxies was already pointed out in simulations with stellar particles (e.g. Tarumi et al. 2021; Goater et al. 2024). In theory, post-processing techniques sampling the IMF could be used to transform such stellar particles into individually resolved stars (e.g. Krumholz et al. 2015; Grand et al. 2018; Sanderson et al. 2020; Gray et al. 2024) and test observational pipelines. However, up-sampling techniques can create strong phase-space artifacts, especially when stellar sampling is limited (see e.g. discussion in Lim et al. 2024). The vastly improved stellar sampling of our resolved-star simulations – $\approx 100,000$ stars per galaxy compared to $\approx 500$ stellar particles in (Rey et al., 2019) – significantly improves the robustness in predicting the kinematic and chemical properties of the stellar distribution. We turn next to quantifying how this impacts the observational recovery of structural parameters and will explore radial trends and properties of distant members in future work.
@@ 5 Discussion and Conclusion
We acknowledge Mordecai-Mark Mac Low for helpful comments improving the quality of this manuscript. EA acknowledges support from NASA ATP grant 80NSSC24K0935 and NSF grant AST23-07950. MR is supported by the Beecroft Fellowship funded by Adrian Beecroft. OA acknowledges support from the Knut and Alice Wallenberg Foundation, the Swedish Research Council (grant 2019-04659), and the Swedish National Space Agency (SNSA Dnr 2023-00164). AP was supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 818085 GMGalaxies). JIR would like to acknowledge support from STFC grants ST/Y002865/1 and ST/Y002857/1. This work used the DiRAC Data Intensive service (DIaL2 / DIaL) at the University of Leicester, managed by the University of Leicester Research Computing Service on behalf of the STFC DiRAC HPC Facility (www.dirac.ac.uk). The DiRAC service at Leicester was funded by BEIS, UKRI and STFC capital funding and STFC operations grants. DiRAC is part of the UKRI Digital Research Infrastructure. We thank the Astrophysics Data Service funded by NASA under Cooperative Agreement 80NSSC21M00561 and the arXiv preprint repository for providing services that were used extensively in this work.

### v2

@@ 1 Introduction
In recent years, numerical models of low-mass galaxies have improved significantly (see, e.g., Hu et al., 2016; Wheeler et al., 2019; Agertz et al., 2020; Smith, 2021), even unlocking the ability to account for individual stars within a whole galaxy (e.g., Hu et al. 2017; Emerick et al. 2018; Andersson et al. 2020, 2021, 2023; Lahén et al. 2020, 2023, 2024a; Calura et al. 2022; Gutcke et al. 2021, 2022a, 2022b; Hislop et al. 2022; Steinwandel et al. 2023, 2024; Deng et al. 2024), as opposed to aggregating stars into a single stellar population particle. By explicitly drawing from the initial mass function (IMF) at stellar birth and modeling the stellar evolution tracks of each star, these models provide a much more detailed and robust account of how, when, and where feedback and metals are deposited into the interstellar medium (ISM). Furthermore, this novel model approach opens up the door for incorporating aspects of collisional dynamics present in star clusters that potentially affect both the ISM, galactic winds, and circumgalactic medium (e.g., runaway stars and clustering of feedback, see Naab & Ostriker, 2017, for a review). While high-order direct $N$-body integration is prohibitively computationally expensive for all stars in a galaxy over a Hubble time, its effects can be treated with sub-grid models (e.g., Andersson et al., 2020, 2023; Steinwandel et al., 2023) using calibrations from highly detailed models applied in idealized set-ups (Grudić et al., 2021, 2022; Lahén et al., 2024b; Polak et al., 2024).
@@
In the rare cases when individual-star models have included a full cosmological environment (e.g. Gutcke et al. 2022a, b; Calura et al. 2022), only massive (and thus less numerous) stars ($\gtrapprox 4\,\,{\rm M}_{\odot}$) are individually sampled to ease the computational load. But slow-moving stellar mass loss from low-mass evolved stars has characteristic velocities ($\approx 10\,\rm km\,s^{-1}$; Höfner & Olofsson 2018) comparable to the depth of gravitational potential wells of faint dwarf galaxies. As a result, winds from e.g. asymptotic giant branch stars (AGBs) likely play a role in chemical enrichment and possibly delay the restart of star formation in galaxies quenched by reionization (e.g. Rey et al. 2020). Furthermore, the lowest mass galaxies are star formation quenched during the epoch of reionization and hence their mass budget in massive stars by redshift zero is fully depleted. Instead dwarf galaxy properties are characterized through the imaging and spectroscopy of the light from their old, evolved low-mass stars, motivating a complete account of the stellar population.
In this paper, we combine all these advances for the first time to take a major step toward interpreting resolved-star data from dwarf galaxies. We perform high-resolution cosmological simulations within the EDGE [^1] Engineering Dwarfs at Galaxy formation’s Edge framework, combining them with the star-by-star INFERNO [^2] INdividual stars with Feedback, Enrichment, and Realistic
Natal mOtions
model (Andersson et al., 2023). With this, we achieve a high numerical resolution ($3\,{\rm pc}$ at $z=0$) that allows us to capture the evolution of supernova (SN) explosions self-consistently over the full cosmological formation of the dwarf, while individually sampling all stars down to $0.5\,\,{\rm M}_{\odot}$ (main sequence lifetime $>14\,\rm Gyr$). This ensures that all major sources of stellar feedback are accounted for on a star by star basis, while self-consistently populating the observable spatial and color-magnitude distribution at $z=0$, at which point stars born before reionization and with mass $\sim 0.8\,{\rm M}_{\odot}$ evolve to the giant branch (thus become bright enough to be observable).
@@ 2 Star-by-star cosmological simulations
Star formation follows Agertz et al. (2020), assuming stars are born in quantities of $500\,{\rm M}_{\odot}$ in cells with gas density $>300\,{\rm cm}^{-3}$ and temperature $<10^{3}\,{\rm K}$. Spawning stars in groups of $500\,{\rm M}_{\odot}$ minimizes issues related to incomplete sampling of the IMF (Smith, 2021). In addition to the gas velocity, stars are stirred with a velocity dispersion of $0.01\,{\rm km\,s}^{-1}$ at birth (Andersson et al., 2023). This velocity dispersion is below the resolution of the gravitational softening (cell size) and serves as a numerical approximation to allow particles with identical initial position end up on trajectories that eventually diverge due to small but repeated dynamical perturbation.
@@
Stellar feedback in our models includes SNe (core-collapse and Ia) and stellar winds (from OB type and AGB stars) implemented as in Andersson et al. (2023). While winds are a sub-dominant component of the feedback budget (in particular at low metallicity, see, e.g., Vink et al., 2001; Mokiem et al., 2007; Vink, 2011), they are an important contributor to metal enrichment and can provides a prompt source of energy injection affecting local star formation (Andersson et al., 2024). Radiation is not included in our suite of simulations (or in the simulations from Rey et al. 2019 we compare to) and is expected to further regulate star formation, systematically decreasing the total stellar masses of our galaxies (Agertz et al. 2020, Rey et al. in prep). We read the chemical yields and mass-loss of each star as a function of mass and metallicity using tabulated values from NuGrid (Pignatari et al., 2016; Ritter et al., 2018) for core-collapse SNe and stellar winds, and from Seitenzahl et al. (2013) for SNe Ia. In all cases, we calculate the main sequence lifetime of a star following the fitting function from Raiteri et al. (1996). SNe Ia follow the time-delay distribution for field galaxies from Maoz & Graur (2017) assuming a rate of $I_{\rm Ia}=2.6\times 10^{-13}\,{\rm yr}^{-1}\,{\rm M}_{\odot}^{-1}$ starting $38\,{\rm Myr}$s after birth.
As in Andersson et al. (2023), massive stars ($8\leq m_{\star}\leq 60\,{\rm M}_{\odot}$) inject fast ($1000\,{\rm km\,s}^{-1}$) winds while on the main sequence and end their life with core-collapse directly injecting material and $10^{51}\,{\rm erg}$ if $m_{\star}\leq 30\,{\rm M}_{\odot}$ (assuming direct collapse to black hole otherwise). While the high resolution of our simulations allows direct injection of SNe energy to obtain accurate capture of the final momentum (Ohlin et al., 2019) in almost all cases ($>96\%$), a switch to terminal momentum injection (Kim & Ostriker, 2015) is applied in the rare cases when the resolution is not sufficient (see Agertz et al., 2020; Andersson et al., 2023, for further details).
@@ 3 Stellar mass growth
Compared with the traditional stellar-particle approach (dotted; Rey et al. 2019), swapping to individual-star feedback systematically increases $M_{\star}$ at all redshifts, by at most a factor of 2. Such variations are well within theoretical uncertainties at this mass scale (see, e.g., the discussion in Agertz et al. 2020, and review in Sales et al. 2022) which should be expected given the extensive updates to the chemical enrichment channels (and thus gas cooling) and updates to the stellar feedback implementation. [^3] Even if the total feedback budget and injection mechanisms are fixed, the location and clustering of individual stars can be significantly different and affect galactic outflows (e.g. Andersson et al. 2020, 2023; Smith 2021).
@@ 4.1 Color-magnitude diagrams
All our dwarf galaxies have diffuse light profiles, with hierarchical merging leading to distant bound members up to $\geq 5\,r_{h}$ similar to those reported in recent observations (e.g., Chiti et al. 2021; Sestito et al. 2023). As we will see, these multi-component spatial distributions can bias the recovery of faint dwarf structural parameters, particularly when few stars are bright enough to be observed. Such outer stellar halo for faint dwarf galaxies was already pointed out in simulations with stellar particles (e.g. Tarumi et al. 2021; Goater et al. 2024). In theory, post-processing techniques sampling the IMF could be used to transform such stellar particles into individually resolved stars (e.g. Krumholz et al. 2015; Grand et al. 2018; Sanderson et al. 2020; Gray et al. 2024) and test observational pipelines. However, up-sampling techniques can create strong phase-space artifacts, especially when stellar sampling is limited (see e.g. discussion in Lim et al. 2024). The vastly improved stellar sampling of our resolved-star simulations – $\approx 100,000$ stars per galaxy compared to $\approx 500$ stellar particles in (Rey et al., 2019) – significantly improves the robustness in predicting the kinematic and chemical properties stellar distribution. We turn next to quantifying how this impacts the observational recovery of structural parameters and will explore radial trends and properties of distant members in future work.
@@ 5 Discussion and Conclusion
We thank the anonymous referee for comments and suggestions that greatly improved the quality of this work. We acknowledge Mordecai-Mark Mac Low for helpful comments improving the quality of this manuscript. EA acknowledges support from NASA ATP grant 80NSSC24K0935 and NSF grant AST23-07950. MR is supported by the Beecroft Fellowship funded by Adrian Beecroft. OA acknowledges support from the Knut and Alice Wallenberg Foundation, the Swedish Research Council (grant 2019-04659), and the Swedish National Space Agency (SNSA Dnr 2023-00164). AP was supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 818085 GMGalaxies). JIR would like to acknowledge support from STFC grants ST/Y002865/1 and ST/Y002857/1. This work used the DiRAC Data Intensive service (DIaL2 / DIaL) at the University of Leicester, managed by the University of Leicester Research Computing Service on behalf of the STFC DiRAC HPC Facility (www.dirac.ac.uk). The DiRAC service at Leicester was funded by BEIS, UKRI and STFC capital funding and STFC operations grants. DiRAC is part of the UKRI Digital Research Infrastructure. We thank the Astrophysics Data Service funded by NASA under Cooperative Agreement 80NSSC21M00561 and the arXiv preprint repository for providing services that were used extensively in this work.